OpenStudy (anonymous):

find the derivative F(x)=(1/(x-f(x)))^4 ?

OpenStudy (jonnyvonny):

Not that I don't want to help you, but this question is a pain the the buttchecks. Just remember that f'(x) = F(x), if I remember correctly. If its not, then leave f'(x) as is. Also, you're doing the chain rule into a quotient rule.

OpenStudy (aum):

\[ F(x)= \left( \frac{1}{ x-f(x) }\right)^4 \\ F'(x) = 4 * \left(\frac{1}{ x-f(x) }\right)^3 * \frac{d}{dx} \left(\frac{1}{ x-f(x) }\right)= ? \]